Shift Summary 27 June 1999, Day

Nickolas Solomey (solomey@hep.uchicago.edu)
Sun, 27 Jun 1999 15:16:58 -0500

Shift Summary for Day shift, 06/27/99

On Shift: Nick and Earl

Others present: Ed (by phone which was often enought to mention), Bob T,
Dave N., Jason H. (ie Jason2), and Hogan

Hours of beam delivered at all intensities: 8.25
Hours of beam delivered at nominal intensity: 8.25 (5e12)
Hours of E832 data (special runs not included): 8.25
Hours of special runs (laser scans, muon runs...): 0 (no special runs
scheduled for this shift)

Accounting of lost time: 10 min recovery from vacuum problems, plus a
handfull of missing spills.

========================================================================
Runs ended or started this shift: None
13745 ongoing from previous run and continues into the next shift.

Tapes brought to FCC this shift: None
(including tapes for online split)

========================================================================
Accesses during shift: None

========================================================================
Description of Shift:
>7:00 Arrived on shift and found the DAQ and CsI all in good order,
but no beam. Estimate 30 min for vacuum problems.
>7:10 Beam returns.
>8:15 finished scanning reference plots, no visible problems with any
detector elements except HA last paddle which was know from previous
shifts.
>9:15 Accelertor missed two spills, but then came back to previous
intensity 4.5e12.
>9:45 Slow DAQ kept having Epicure errors, noticed that during the
spill the magnet nm1u which is set to 1670 and read back at 1630 to 1650
will ocassionally drop in the spill to 1440! Called the MCR to
point this out to them, they said he suspected it was not a problem
since it does not effect the beam, but he will keep an eye on it.
>10:05 Paged Rick Ford to describe to him the nm1u magnet problem which
is visibly linked to a shift in h of the beam location when the magnet
drops during the spill for one scan by 200 amps.
>10:50 Lost two spills due to having to run Kfrend.
>12:20 Finished another scan threw the online plots all is fine
except in the hotbits plot for the HCC the upper left quandrant is
all black this is an old problem that Sydney knows about, there
is also a fitting problem in the lowest energy curve for the etot plot
but the data looks good. Experts should look into fixing these when
there is time permitting.
>12:40 Found REG/SEM = 0.582 E5 at spill 293 -- increasing after that.
>13:30 Both check lists were done.
>14:15 Finished another scan threw online plots. This time in dc#summary
plane 15 of the DC effiencies had a error, but everything looked like
the sample graph so it was not clear why there was the unhappy-face,
Ed is looking into it.
>14:22 Ed calls back with followup on the DC plane 15 problems and
suggests we page Hogan and have him look at it from home. The idea is
that if there is an unhappy-face and it is normal then someone should
change the sensativity level, the author of this summary graph should
look into it when they get in monday.
>14:30 Hogan arrives, says DC plane 15 unhappy because of a lot of
cells on edge below 97% effiency and there is not much he can do
about that so we will live with it.

=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=
Dr. Nickolas Solomey, Research Scientist, Enrico Fermi Institute
The University of Chicago, 5640 S. Ellis Ave, Chicago, Ill. 60637 USA
Phone: (773) 702-3278, fax: x-6317, http://hep.uchicago.edu/~solomey